OpenStudy (anonymous):

Using the quadratic formula how would one solve 5x^2 -4px-p^2= 0 [i'm really confused by the two powers; shouldn't there only be one in a quadratic pattern]

OpenStudy (anonymous):

\[5x^2-4px-p^2\]

OpenStudy (anonymous):

\[b ^{2}-4ac=0\] a=5, b=-4p, c=-p^2

OpenStudy (anonymous):

use quadratic formula to solve the p

OpenStudy (anonymous):

So you solve for p and x seperatley

OpenStudy (anonymous):

the question want to find p right?

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

use factorisation (-5x-p)(-1x+p)=0

OpenStudy (anonymous):

here you are required to factorise to solve for x ..good job @jckm135

OpenStudy (anonymous):

But it says use quadratic formula

OpenStudy (anonymous):

honestly i've never used the quadratic formula for this type of example

OpenStudy (anonymous):

ill try wolfram alpha, thanks anyway

OpenStudy (helder_edwin):

when u have an expression like this u have to ask which of the two letters is your unknown

OpenStudy (anonymous):

if anyone is interested http://i.imgur.com/GsybwCP.png

ganeshie8 (ganeshie8):

p is constant, x is variable here

OpenStudy (helder_edwin):

if the unknown is x then your equation is \[ \large \color{red}{5}x^2\color{red}{-4p}x\color{red}{-p^2}=0 \] and the quadratic formula gives you the solutions \[ \large x=\frac{-(-4p)\pm\sqrt{(-4p)^2-4(5)(-p^2)}}{2(5)} \]

OpenStudy (helder_edwin):

got it? and yes, as ganeshie8 points out, it is traditional for x to be the unknown and any other letter to be a constant

OpenStudy (anonymous):

Yeah thanks a lot i think i just got a bit lost when i was adding the like terms